You are performing a two-tailed test. If α = .004 , find the positive critical value, to three decimal places. zα/2 =

You are performing a left-tailed z-test

If α=.025, and your test statistic is z=−1.75, do you:

  • Reject Null Hypothesis
  • Fail to Reject Null Hypothesis
Answer #1

Solution:

= 0.004

/ 2 = 0.004 / 2 = 0.002

Z/2 = Z0.002 = 2.878

positive critical value = 2.878

= 0.025

z = -1.75

P(z < -1.75) = 0.0401

P-value = 2 * 0.0401 = 0.0802

P-value >

Fail to Reject Null Hypothesis